OpenStudy (mrs.biersackxx):

A cylindrical well is 20 meters deep and has a diameter of 1.5 meters. Approximately how many cubic meters of soil were dug out to make the well? (Use π = 3.14.)

Vocaloid (vocaloid):

first find the radius

OpenStudy (mrs.biersackxx):

so is that half of 1.5?

Vocaloid (vocaloid):

yup! that gives us 0.75

Vocaloid (vocaloid):

then we use the formula v = (pi)(r^2)(h) where r = 0.75 and h = 20

OpenStudy (mrs.biersackxx):

ah ok

Vocaloid (vocaloid):

and pi = 3.14

OpenStudy (mrs.biersackxx):

so then u multiply all of it?

Vocaloid (vocaloid):

yeah

OpenStudy (mrs.biersackxx):

i got 47.1

OpenStudy (mrs.biersackxx):

i think its wrong

Vocaloid (vocaloid):

nope, try again

Vocaloid (vocaloid):

v = (3.14)*(20)*(.75)^2 = ?

OpenStudy (mrs.biersackxx):

Oh i think i did something wrong

OpenStudy (mrs.biersackxx):

agh idk what im doing wrong

Vocaloid (vocaloid):

just multiply from left to right

Vocaloid (vocaloid):

(3.14)*20 = ?

OpenStudy (mrs.biersackxx):

62.8

Vocaloid (vocaloid):

good, now multiply that by (0.75)*(0.75)

OpenStudy (mrs.biersackxx):

35.33

OpenStudy (mrs.biersackxx):

OH I GOT IT

Vocaloid (vocaloid):

good, thats your final answer ~

OpenStudy (mrs.biersackxx):

thanks ur simply amazing